Which expression is equivalent to sin θ + cos θ cot θ?
The choices are:
A cot θ
B tan θ
C sec θ
D csc θ
Please give an answer and clear work/explanation!

Answer by Alan3354(69443)

Which expression is equivalent to sin θ + cos θ cot θ?
= sin + cos*(cos/sin)
= sin + cos^2/sin
= (sin^2 + cos^2)/sin
= 1/sin
= csc
=================